1.During 2013, a country reported that its real GDP increased by $3.0 billion. If the slope of its aggregate planned expenditur

2.During 2013, exports increase from $1.0 trillion to $1.5 trillion. If the slope of the aggregate planned expenditure (AE) cur
Answer #1

Q1
Answer
Multiplier =1/(1-MPC)
=1/(1-0.9)
=10
change in GDP is because of change in one of the factors of the AD
it increases in one or some or all of the Exports, government spending, investment, consumption
By
=change in GDP/multiplier
=3/10
=0.3 billion
the change in 0.3 billion of increase is option 1 where export increases by $0.3 billion.
=====
Q2
Answer
Multiplier =1/(1-MPC)
=1/(1-0.75)
=4

change in GDP=change in export * multiplier
=(1.5-1)*4
=$2 trillion

option 1.